# Meta-Learning Reinforcement Learning for Crypto-Return Prediction

### v1

@@ Abstract
Junqiao Wang1,2, Zhaoyang Guan3, Guanyu Liu10, Tianze Xia4, Xianzhi Li5, Shuo Yin8, Xinyuan Song9 Chuhan Cheng1, Tianyu Shi6, Alex Lee7, 1Sichuan University, 2Shanghai Artificial Intelligence Laboratory, 3Northwestern University, 4Huazhong University of Science and Technology, 5Queen’s University, 6University of Toronto, 7TrueNorth, 8Tsinghua University, 9Emory University, 10University of Macau Correspondence: ty.shi@mail.utoronto.ca
@@ 1 Introduction
To address the challenges in cryptocurrency prediction, we propose Meta-RL-Crypto, a framework that combines meta-reward-driven self-improvement with multi-modal trading intelligence in Figure LABEL:fig:architecture. The central innovation is a triple-loop learning process in which a single large language model (LLM) takes on three distinct roles. First, the Actor processes on-chain metrics (such as gas fees and transaction graphs), news, and sentiment to generate next-day forecasts for crypto-assets Meister and Price (2024); Liu and Jia (2025). Next, the Judge evaluates these forecasts using a multi-objective reward vector, which incorporates absolute returns, the Sharpe ratio, drawdown control, and sentiment alignment. This approach reduces potential bias from relying on a single metric. Finally, the Meta-Judge refines the Judge’s reward policy through preference comparisons, helping to prevent reward drift and length bias Lee et al. (2024); Wu et al. (2024). This closed-loop system enables continuous self-improvement without human intervention, allowing it to adapt dynamically to shifts in the market.
@@ 2.1 LLMs for Financial Analysis and Cryptocurrency Markets
Recent advancements in large language models (LLMs) have transformed financial-market research, though most studies have concentrated on equities, leaving the rich on-chain data of cryptocurrencies less explored Roumeliotis et al. (2024); Li et al. (2024). Early time-series forecasting in finance evolved from traditional econometrics to machine learning and reinforcement learning methods, with Long Short-Term Memory networks proving effective for sequential price data Fjellström (2022); Siami-Namini and Siami Namin (2018); Sezer et al. (2019). Transformer-based models, such as Informer Zhou et al. (2021), AutoFormer Wu et al. (2021), PatchTST Nie et al. (2022), and TimesNet Liu et al. (2023b), now lead the field in long-horizon forecasting. In addition, domain-specific LLMs—such as FinGPT Liu et al. (2023a), BloombergGPT Wu et al. (2023), and FinMA Xie et al. (2023)—expand LLM capabilities to sentiment analysis, entity recognition, question answering, and market prediction. While these techniques show promise for more robust decision-making, their application to cryptocurrency markets remains underexplored. Our work addresses this gap by integrating both on-chain and off-chain signals with self-reflective LLM agents to navigate the volatile, information-rich landscape of cryptocurrency assets.

### v2

@@ Abstract
Junqiao Wang* 1, Zhaoyang Guan* 2, Chuhan Cheng* 1, Guanyu Liu9, Tianze Xia3, Xianzhi Li4, Shuo Yin7, Xinyuan Song8, Tianyu Shi5, Alex Lee6, 1Sichuan University, 2Shanghai Artificial Intelligence Laboratory, 3Northwestern University, 4Huazhong University of Science and Technology, 5Queen’s University, 6University of Toronto, 7TrueNorth, 8Tsinghua University, 9Emory University, 10University of Macau Correspondence: ty.shi@mail.utoronto.ca
@@ 1 Introduction
To address the challenges in cryptocurrency prediction, we propose Meta-RL-Crypto, a framework that combines meta-reward-driven self-improvement with multi-modal trading intelligence in Figure 1. The central innovation is a triple-loop learning process in which a single large language model (LLM) takes on three distinct roles. First, the Actor processes on-chain metrics (such as gas fees and transaction graphs), news, and sentiment to generate next-day forecasts for crypto-assets Meister and Price (2024); Liu and Jia (2025). Next, the Judge evaluates these forecasts using a multi-objective reward vector, which incorporates absolute returns, the Sharpe ratio, drawdown control, and sentiment alignment. This approach reduces potential bias from relying on a single metric. Finally, the Meta-Judge refines the Judge’s reward policy through preference comparisons, helping to prevent reward drift and length bias Lee et al. (2024); Wu et al. (2024). This closed-loop system enables continuous self-improvement without human intervention, allowing it to adapt dynamically to shifts in the market.
@@ 2.1 LLMs for Financial Analysis and Cryptocurrency Markets
Recent advancements in large language models (LLMs) have transformed financial-market research, though most studies have concentrated on equities, leaving the rich on-chain data of cryptocurrencies less explored Roumeliotis et al. (2024); Li et al. (2024). Early time-series forecasting in finance evolved from traditional econometrics to machine learning and reinforcement learning methods, with Long Short-Term Memory networks proving effective for sequential price data Fjellström (2022); Siami‐Namini and Siami Namin (2018); Sezer et al. (2019). Transformer-based models, such as Informer Zhou et al. (2021), AutoFormer Wu et al. (2021), PatchTST Nie et al. (2022), and TimesNet Liu et al. (2023b), now lead the field in long-horizon forecasting. In addition, domain-specific LLMs—such as FinGPT Liu et al. (2023a), BloombergGPT Wu et al. (2023), and FinMA Xie et al. (2023)—expand LLM capabilities to sentiment analysis, entity recognition, question answering, and market prediction. While these techniques show promise for more robust decision-making, their application to cryptocurrency markets remains underexplored. Our work addresses this gap by integrating both on-chain and off-chain signals with self-reflective LLM agents to navigate the volatile, information-rich landscape of cryptocurrency assets.